If a vector 3i+4j is rotated by 90 degree then it is changed by;

Grade:12th pass

1 Answers

Khimraj
6 years ago
let new vector is ai + bj.
since magnitude will remains same
so 
a+ b2 = 25   – (1)
and it will perpendicular to 3i+4j. So dot product will be zero.
3a + 4b =0
a = -4b/3.
put this in equation (1)
16b2/9 + b2 = 25
b = +3, -3.
So a = -4, +4.
So new vector will be -4i + 3j or 4i – 3j.
Hope it clears.
If you like answer then please approve it.